OpenStudy (anonymous):

which are periodic and which are not ?how they are periodic? [cos(2t-i/3)]^2. x[n]=cos(pi/8*n^2).

OpenStudy (anonymous):

If you can prove \(\mathsf{f(x) =f(x+t)}\), where \(\mathsf t\) is a constant, then the function \(\mathsf{f(x)}\) is Periodic. In your first problem, \(\large \mathsf {f(x)=\left[\cos\left(2t-\frac{i}{3}\right) \right]^2 }\). Before I begin I have few questions. What is '\(\mathsf i\)' ? And is '[' and ']' for floor function?

OpenStudy (anonymous):

\[\cos(x)\] testing one two

OpenStudy (anonymous):

@ Ishaan94 'i' is periodic and the '[]' is for discrete time signal.